1. Water is flowing at a speed of 2ms^-1 in a pipe of cross sectional area 0.02m^2 at point 1. If the cross sectional area is reduced to half at point 2, find the velocity at point 2.
1
Expert's answer
2019-08-16T09:19:06-0400

Continuity equation states


"A_1v_1=A_2v_2"

So


"v_2=\frac{A_1}{A_2}v_1=\frac{A_1}{A_1/2}\times 2\:\rm{m/s}=4\:\rm{m/s}"